# Sensitivity of the Global 21-cm Signal to Dark Matter-Baryon Scattering

### v1

@@ Abstract
Keywords
Cosmology, Dark matter, 21-cm global signal, Dark ages, Reionization
preprint: APS/123-QED
I Introduction
Many observational probes testify to the existence of Dark Matter (DM), which comprises around $\sim$ 85% of the matter content in the universe at present [1, 2, 3, 4, 5]. Despite numerous lines of evidence for the existence of DM, its fundamental nature still remains elusive. In standard cosmology, DM is typically modeled as a cold, non-relativistic, collisionless particle, often referred to as cold dark matter (CDM). However, CDM may not be able to account for the full range of observational anomalies reported on various scales [6, 7, 8, 9, 10, 6, 7, 10, 11, 12], which highlights the importance of exploring alternative models beyond CDM. We focus on interacting DM (IDM) [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53], which features elastic scattering with the Standard Model particles. Significant efforts have been made to search for signatures of IDM in data from a variety of probes [54, 55], such as direct detection experiments of DM [29, 56, 57, 58, 59], cosmic rays [60, 61], cosmic microwave background (CMB) observations [27, 22, 15, 28, 29], the Lyman-$\alpha$ forest flux [16, 15, 17], and Milky Way satellite abundance [23, 28]. Complimentary to these experiments, observations of the 21-cm signal from the high-redshift universe can also provide unique exploration opportunities [62, 63, 64, 65, 35, 36]. IDM models, especially those with Coulomb-like interactions, have garnered significant attention due to their potential to alter the global 21-cm signal from neutral hydrogen at cosmic dawn. Motivated by the distinct effect of IDM on the global 21-cm signal, we focus this study on quantifying the potential of 21-cm cosmology as a probe of IDM.
@@ I Introduction
This study explores the sensitivity of the global 21-cm experiments to detecting the effects of DM interactions. Following previous literature [16, 28, 31, 33, 32, 13], we consider IDM models in which the interaction cross section is parameterized as $\sigma(v)=\sigma_{\mathrm{0}}v^{n}$, where $\sigma_{0}$ is the unknown cross section normalization and $v$ is the relative velocity between DM and baryons. We focus on two models: i) a Coulomb-like interaction ($n=-4$[18, 19, 20, 40, 53] with $n=-4$, and ii) a velocity-independent interaction with $n=0$ [21, 22, 23, 24, 25, 26]. To model the global 21-cm signal in presence of DM-baryon scattering, we follow the approach of Ref. [40], and include the effects of scattering on structure formation and on the thermal history of baryons, using the semi-analytic merger-tree code Galacticus [87], and the 21-cm signal calculator Ares [88]. Using the Fisher matrix formalism, we analyze four distinct experimental scenarios and forecast the sensitivity to recovering the IDM cross section and astrophysical parameters, exploring the degeneracies between them. We find that for the $n=0$ case, a SARAS-like (similar configuration to SARAS3) experiment has a sensitivity comparable to the strongest bounds to date obtained from the Milky Way satellite abundance measurements [23, 25]. Furthermore, other considered experimental scenarios, which include an EDGES-like experiment (similar configuration to EDGES), Future1 (more integration time compared to EDGES), and Future2 (more integration time and frequency channels compared to EDGES), show stronger sensitivities. For $n=-4$ case, all four scenarios show improvements compared to the strongest bounds obtained from the cosmic microwave anisotropy (CMB) measurements [28].
This paper is organized as follows. In Section II, we summarize the IDM models and outline our approach for modeling the global 21-cm signal. In Section III, we summarize the Fisher matrix formalism used to calculate the forecasts for various experiment scenarios. In Section IV, we present the results and discuss their implications. We summarize the main findings of this in Section V and briefly discuss the future directions.
@@
II Global 21-cm Signal with IDM
@@
We investigate a specific IDM model where DM interacts with baryons through elastic scattering, resulting in the momentum and heat transfer between the corresponding cosmological fluids, and a dissipation of the relative bulk velocity. For non-relativistic scattering, the momentum-transfer cross section can be parameterized as $\sigma(v)=\sigma_{\mathrm{0}}v^{n}$, where $v$ is the relative velocity between the interacting particles and $\sigma_{\mathrm{0}}$ is the free fitting parameter of the model, representing the cross section normalization. We consider two scenarios, a Coulomb-like interaction ($n=-4$) and velocity-independent interaction ($n=0$). With Coulomb-like interaction, DM is assumed to scatter with both charged particles and neutral targets (hydrogen and helium atoms), following Refs. [38, 50, 40].
@@
To model the effects of IDM on cosmology, we use a modified version of the publicly available code class [^1] https://github.com/kboddy/class_public/tree/dmeff [28, 22, 21] to obtain the linear matter power spectrum, evolution of DM and baryon temperatures, and the relative bulk velocity, in presence of interactions. As noted in previous literature, in case of a Coulomb-like interaction, the matter power spectrum is affected across a range of scales, with a characteristic scale-dependent suppression, with an onset at $k\approx 10^{-2}\ h\ \rm{Mpc}^{-1}$; on smaller scales, the transfer function shows a decrement of power (see Figure $3$ in Ref. [40]), with the suppression inversely proportional to $\sigma_{\mathrm{0}}$. In the velocity-independent scattering model, the suppression of the matter power spectrum is similar to that of warm DM, and the interactions in the late-time are negligible [25].
Figure 1: The Global 21-cm signal as a function of frequency in IDM models with $n=0$ (left-hand panel) and $n=-4$ (right-hand panel), for DM particle mass of $1$ GeV, are shown. The error bars represents the rms noise of the EDGES experiment. The color scheme captures the change in the signal as a function of the interaction cross section.
@@
$\frac{dn}{dM}=f(\sigma)\frac{\rho_{\mathrm{m}}}{M}\frac{d\ln(\sigma^{-1})}{dM}\,,$ (1)
@@
$\sigma^{2}(M)=\frac{1}{2\pi^{2}}\int^{\infty}_{\mathrm{0}}4\pi k^{2}P(k)W^{2}(k|M)\mathrm{d}k\,.$ (2)
@@
$k_{\mathrm{s}}(M)=2.5/R\quad\text{and}\quad R=\left(\frac{3M}{4\pi\bar{\rho}}\right)^{1/3}.$ (3)
@@
$f(\sigma)=A\left[\left(\frac{\sigma}{b}\right)^{-a}+1\right]e^{-c/\sigma^{2}},$ (4)
where $A$ controls the amplitude, $a$ controls the tilt, $b$ sets the mass scale at which the power law becomes significant, and $c$ determines the high-mass cutoff scale above which halo abundances exponentially decrease. This function is widely used, and it has successfully reproduced the HMFs at low z predicted by N-body simulations [94, 95]. We use galacticus [^2] https://github.com/galacticusorg/galacticus code, which utilizes the ePS formalism, to calculate the HMF for the given initial conditions generated using CLASS up to $z=500$. We choose $10^{5}\mathrm{M}_{\odot}$ as the resolution in galacticus in order to capture the details of the smallest scales important for the Global signal. As discussed in Ref. [40], the size of the sound horizon is evolving, and due to interactions, the structure formation may be affected, especially at late times. Such interactions can suppress the growth of small-scale structures, leading to a modification in the HMF by reducing the abundance of low-mass halos compared to the standard CDM scenario.
II.1 21-cm Global Signal
@@
$\delta T_{\mathrm{b}}\simeq$ $27(1-\bar{x}_{i})\left(\frac{1-Y_{p}}{0.76}\right)\left(\frac{\Omega_{b}h^{2}}{0.023}\right)\left(\frac{0.15}{\Omega_{m}h^{2}}\frac{1+z}{10}\right)^{1/2}$
×$\times\left(1-\frac{T_{\mathrm{\gamma}}}{T_{\mathrm{S}}}\right)mK,$ (5)
where $x_{i}$ is the mean ionized fraction, $Y_{\mathrm{p}}$ is the Helium mass fraction, $T_{S}$ is the spin temperature, and $T_{\mathrm{\gamma}}$ is the CMB temperature. In this work, we use ares [^3] https://github.com/treydriskell/ares/tree/galacticus_hmf [88] to predict the Global signal utilizing the output obtained from galacticus for a given set of astrophysical and IDM parameters. In the computation of the global 21-cm signal, ares takes into account for the Lyman-$\alpha$ flux, X-ray heating, and ionizing emission effects.
The spin temperature, $T_{\mathrm{S}}$, in Eq. II.1 depends on three competing processes: coupling to the CMB photons, collisional coupling with the gas, and Lyman-$\alpha$ coupling:
$T_{\mathrm{S}}^{-1}=\frac{T_{\mathrm{\gamma}}^{-1}+x_{\mathrm{c}}T_{\mathrm{K}}^{-1}+x_{\mathrm{\alpha}}T_{\mathrm{\alpha}}^{-1}}{1+x_{\mathrm{c}}+x_{\mathrm{\alpha}}},$ (6)
where $x_{\mathrm{c}}$ is the collisional coupling coefficient [97], $T_{\mathrm{K}}$ is the kinetic temperature of the gas, and $x_{\mathrm{\alpha}}$ is the coefficient of Lyman-$\alpha$ coupling with $T_{\mathrm{\alpha}}$ being the corresponding Lyman-$\alpha$ color temperature.
@@
$\epsilon_{\nu}=\mathrm{f}_{\ast}\overline{\rho}_{\mathrm{b,0}}\frac{\mathrm{d}\mathrm{f}_{\rm{coll}}}{\mathrm{dt}}\mathrm{l}_{\nu},$ (7)
@@
where $\mathrm{f}_{\ast}$ is the star formation efficiency, $\overline{\rho}_{b,0}$ is the mean baryon density, $\mathrm{f}_{\rm{coll}}$ is the fraction of matter in collapsed halos with virial temperatures in excess of $\mathrm{T}_{\mathrm{min}}$, and $l_{\nu}$ encodes the efficiency of photon production (per stellar baryon) as a function of photon frequency. Note that the product $\mathrm{f}_{\ast}\overline{\rho}_{\mathrm{b,0}}\mathrm{df}_{\rm{coll}}/\mathrm{dt}$ is equivalent to the star formation rate density (SFRD) of the Universe in this model.
@@
We treat the Lyman-$\alpha$, Lyman-continuum, and X-ray bands separately, assigning each its own distinct efficiency factor $\mathrm{l}_{\nu}$. For the Lyman-$\alpha$ photons, we set $\mathrm{l}_{\nu}=\mathrm{N}_{\rm{lw}}$, where $\mathrm{N}_{\rm{lw}}$ is the number of photons emitted per stellar baryon. For ionizing photons, $\mathrm{l}_{\nu}=\mathrm{f}_{\mathrm{esc}}\mathrm{N}_{\mathrm{ion}}$. Here, we account for the possibility that only a fraction, $\mathrm{f}_{\mathrm{esc}}$, of the ionizing photons emitted by stars, $\mathrm{N}_{\mathrm{ion}}$, actually escape their host galaxies. Finally, for X-rays, $\mathrm{l}_{\nu}$ is linked to the efficiency of X-ray production in nearby star-forming galaxies. The soft X-ray luminosity $\mathrm{L}_{\mathrm{X}}$ follows the relation $\mathrm{L}_{\mathrm{X}}\simeq 2.6\times 10^{39}\mathrm{erg}\ \rm{s}^{-1}(\mathrm{M}_{\odot}/\rm{yr})^{-1}\left(\mathrm{SFR}/(\mathrm{M}_{\odot}/\rm{yr})\right)$, where SFR is the star formation rate [98]. We introduce a free parameter $\mathrm{f}_{\mathrm{X}}$ that scales this relation, i.e., $\mathrm{f}_{\mathrm{X}}=1$ indicates the local $\mathrm{L}_{\mathrm{X}}/\rm{SFR}$ relation, while $\mathrm{f}_{\mathrm{X}}>1$ indicates more efficient X-ray production (per unit star formation) than $\mathrm{z}\sim 0$ galaxies. Note that $\mathrm{N}_{\mathrm{lw}}$, the product $\mathrm{f}_{\mathrm{esc}}\mathrm{N}_{\mathrm{ion}}$, and $\mathrm{f}_{\mathrm{X}}$ are all degenerate with $\mathrm{f}_{\ast}$. Consequently, parameter forecasts are presented in terms of the product of $\mathrm{f}_{\ast}$ and these parameters, and $\mathrm{f}_{\ast}$ is not independently varied in this work. It is possible to break this degeneracy by independently constraining $\mathrm{f}_{\ast}$ via high-$\mathrm{z}$ galaxy luminosity functions [99], and performing a joint LF/21-cm likelihood analysis [100, 101]. However, we defer an exploration of this approach to future work.
@@
With the UV and X-ray emissivity in hand, one can determine the Lyman-$\alpha$ intensity as a function of redshift, as well as the ionization and heating rates of the intergalactic medium (IGM). These calculations are performed within a two-zone model of the IGM, where the fully-ionized regions and ”bulk IGM” outside ionized bubbles are modeled separately. As a result, the mean ionized fraction $\bar{x}_{i}$ reflects both the volume filling factor of ionized gas, typically denoted as $Q$, and the ionized fraction of the bulk neutral IGM, $x_{e}$, such that $\bar{x}_{i}=Q+x_{e}(1-Q)$. The spin temperature represents the temperature of the largely neutral bulk IGM. For further details, see [88].
@@
We present the 21-cm Global signals computed within the IDM cosmology for various cross section values and $m_{\mathrm{\chi}}=1$ GeV in Fig. 1, along with the error bars for an EDGES-like experiment. The cosmological and astrophysical parameters are kept fixed at their fiducial values for this plot. The left-hand panel corresponds to the $n=0$ (velocity-independent cross section) model, while the right-hand panel represents the $n=-4$ (Coulomb-like interaction) model. From the left-hand panel, it is evident that for the velocity-independent cross section model, the signals show a smooth shift of the troughs towards higher frequencies as the cross section increases. However, in case of Coulomb-like interaction, the signal exhibits a non-monotonic behavior in the position of the troughs with increasing cross sections. This difference in the behavior of the 21-cm signal with cross section for two IDM models can be attributed to the competition of two effects: the dissipation of the relative bulk velocity of DM and baryons and the impact of interaction cross section on the efficiency of baryon cooling.
Experiment $\nu_{\mathrm{min}},\nu_{\mathrm{max}}$ (MHz) $N_{\mathrm{ch}}$ $\Delta\nu$ (kHz) $\tau_{\mathrm{obs}}$ (h) $\mathrm{T}_{\mathrm{rec}}$ (K) $\sigma_{\mathrm{RMS}}\ (\mathrm{mK})$ Input Noise

### v2

@@ Abstract
1 Introduction
Many observational probes testify to the existence of Dark Matter (DM), which comprises around $\sim$ 85% of the matter content in the universe at present [107, 86, 84, 34, 4]. Despite numerous lines of evidence for the existence of DM, its fundamental nature still remains elusive. In standard cosmology, DM is typically modeled as a cold, non-relativistic, collisionless particle, often referred to as cold dark matter (CDM). However, CDM may not be able to account for the full range of observational anomalies reported on various scales [1, 8, 41, 88, 79, 101, 106], which highlights the importance of exploring alternative models beyond CDM. We focus on interacting DM (IDM) [90, 21, 104, 43, 78, 97, 18, 60, 19, 50, 63, 85, 75, 74, 7, 76, 96, 105, 14, 20, 58, 32, 72, 70, 45, 13, 89, 42, 53, 52, 64, 57, 73, 17, 12, 62, 71, 61, 11, 44, 36], which features elastic scattering with the Standard Model particles. Significant efforts have been made to search for signatures of IDM in data from a variety of probes [49, 28], such as direct detection experiments of DM [96, 9, 5, 6, 10], cosmic rays [27, 31], cosmic microwave background (CMB) observations [7, 50, 104, 76, 96], the Lyman-$\alpha$ forest flux [43, 104, 78], and Milky Way satellite abundance [63, 76]. Complimentary to these experiments, observations of the 21-cm signal from the high-redshift universe can also provide unique exploration opportunities [24, 93, 30, 2, 72, 70]. IDM models, especially those with Coulomb-like interactions, have garnered significant attention due to their potential to alter the global 21-cm signal from neutral hydrogen at cosmic dawn. Motivated by the distinct effect of IDM on the global 21-cm signal, we focus this study on quantifying the potential of 21-cm cosmology as a probe of IDM.
@@ 1 Introduction
This study explores the sensitivity of the global 21-cm experiments to detecting the effects of DM interactions. Following previous literature [43, 76, 14, 58, 20, 90], we consider IDM models in which the interaction cross section is parameterized as $\sigma(v)=\sigma_{\mathrm{0}}v^{n}$, where $\sigma_{0}$ is the unknown cross section normalization and $v$ is the relative velocity between DM and baryons. We focus on two models: i) a Coulomb-like interaction (n=4n=-4[97, 18, 60, 42, 36] with n=4n=-4, and ii) a velocity-independent interaction with $n=0$ [19, 50, 63, 85, 75, 74]. To model the global 21-cm signal in presence of DM-baryon scattering, we follow the approach of Ref. [42], and include the effects of scattering on structure formation and on the thermal history of baryons, using the semi-analytic merger-tree code Galacticus [16], and the 21-cm signal calculator Ares [68]. Using the Fisher matrix formalism, we analyze four distinct experimental scenarios and forecast the sensitivity to recovering the IDM cross section and astrophysical parameters, exploring the degeneracies between them. We find that for the $n=0$ case, a SARAS-like (similar configuration to SARAS3) experiment has a sensitivity comparable to the strongest bounds to date obtained from the Milky Way satellite abundance measurements [63, 75]. Furthermore, other considered experimental scenarios, which include an EDGES-like experiment (similar configuration to EDGES), Future1 (more integration time compared to EDGES), and Future2 (more integration time and frequency channels compared to EDGES), show stronger sensitivities. For n=4n=-4 case, all four scenarios show improvements compared to the strongest bounds obtained from the cosmic microwave anisotropy (CMB) measurements [76].
This paper is organized as follows. In Section 2, we summarize the IDM models and outline our approach for modeling the global 21-cm signal. In Section 3, we summarize the Fisher matrix formalism used to calculate the forecasts for various experiment scenarios. In Section 4, we present the results and discuss their implications. We summarize the main findings of this in Section 5 and briefly discuss the future directions.
@@
2 Global 21-cm Signal with IDM
@@
We investigate a specific IDM model where DM interacts with baryons through elastic scattering, resulting in the momentum and heat transfer between the corresponding cosmological fluids, and a dissipation of the relative bulk velocity. For non-relativistic scattering, the momentum-transfer cross section can be parameterized as $\sigma(v)=\sigma_{\mathrm{0}}v^{n}$, where $v$ is the relative velocity between the interacting particles and $\sigma_{\mathrm{0}}$ is the free fitting parameter of the model, representing the cross section normalization. We consider two scenarios, a Coulomb-like interaction (n=4n=-4) and velocity-independent interaction ($n=0$). In both cases, DM particles scatter with neutral hydrogen and helium atoms, following Refs. [13, 61, 42]. To model the Global 21-cm signal in the IDM cosmology, we follow the methodology developed in Ref. [42]. In the following, we briefly summarize this formalism and refer the reader to Ref. [42] for a comprehensive discussion.
The interaction between DM and baryon modifies the evolution of perturbations in the DM-baryon fluid, which introduces some modifications to the linear Boltzmann equations for both species. In synchronous gauge, and assuming both fluids are non-relativistic, the perturbed Boltzmann equations take the following form
$\dot{\delta}_{\chi}$ $=-\theta_{\chi}-\tfrac{\dot{h}}{2},\qquad\dot{\delta}_{b}=-\theta_{b}-\tfrac{\dot{h}}{2},,$
$\dot{\theta}_{\chi}$ $=-\frac{\dot{a}}{a}\,\theta_{\chi}+c_{\chi}^{2}k^{2}\delta_{\chi}+R_{\chi}(\theta_{b}-\theta_{\chi}),$
$\dot{\theta}_{b}$ $=-\frac{\dot{a}}{a}\,\theta_{b}+c_{b}^{2}k^{2}\delta_{b}+\frac{\rho_{\chi}}{\rho_{b}}R_{\chi}(\theta_{\chi}-\theta_{b})+R_{\gamma}(\theta_{\gamma}-\theta_{b}),$ (2.1)
where $\delta_{\chi}$ and $\delta_{b}$ denote the DM and baryon density perturbations, $\theta_{\chi}$ and $\theta_{b}$ are their velocity divergences, $c_{\chi}$ and $c_{b}$ are the respective sound speeds, $h$ is the trace of the scalar metric perturbation, and $R_{\gamma}$ represents the Compton momentum–transfer rate between baryons and photons. Equation (2.1) highlights that the drag terms proportional to $R_{\chi}$ are the point through which the microscopic interaction parameters $(m_{\chi},\sigma_{0})$ influence the evolution of cosmological perturbations. For DM scattering off multiple baryonic targets, the coefficient $R_{\chi}$ is given by
$R_{\chi}\;=\;\sum_{t}N_{n}\,\frac{a\rho_{t}\sigma_{0}}{m_{\chi}+m_{t}}\left(\frac{T_{\chi}}{m_{\chi}}+\frac{T_{K}}{m_{t}}+\frac{V_{\rm rms}^{2}}{3}\right)^{\frac{n+1}{2}},$ (2.2)
@@
where $N_{n}=2^{(n+5)/2}\Gamma\!\left(3+\frac{n}{2}\right)/(3\sqrt{\pi})$ and the sum is over baryonic targets $t$, with mass $m_{t}$ and density $\rho_{t}$, and $T_{\chi}$, $T_{K}$, and $V_{\rm rms}$ denote the DM temperature, baryon temperature, and rms bulk relative velocity, respectively. This expression shows explicitly how the DM mass and cross-section affects the efficiency of momentum and heat transfer between the two fluids. Together, Eqs. (2.1) and (2.2) provide the essential link between the microphysical scattering process and the macroscopic evolution of cosmological perturbations. The DM-baryon interaction also affects the temperature evolution of both the species, thereby influencing the Global 21-cm signal. We use a modified version of the publicly available code class [^1] https://github.com/kboddy/class_public/tree/dmeff [76, 50, 19], which self-consistently computes the linear matter power spectrum, evolution of DM and baryon temperatures, and their relative bulk velocity in the presence of DM-baryon interactions. As noted in previous literature, in case of a Coulomb-like interaction, the matter power spectrum is affected across a range of scales, with a characteristic scale-dependent suppression, with an onset at $k\approx 10^{-2}\ h\ \rm{Mpc}^{-1}$; on smaller scales, the transfer function shows a decrement of power (see Figure $3$ in Ref. [42]), with the suppression inversely proportional to $\sigma_{\mathrm{0}}$. In the velocity-independent scattering model, the suppression of the matter power spectrum is similar to that of warm DM, and the interactions in the late-time are negligible [75].
@@
$\frac{dn}{dM}=f(\sigma)\frac{\rho_{\mathrm{m}}}{M}\frac{d\ln(\sigma^{-1})}{dM}\,$, (2.3)
@@
$\sigma^{2}(M)=\frac{1}{2\pi^{2}}\int^{\infty}_{\mathrm{0}}4\pi k^{2}P(k)W^{2}(k|M)dk\,.$ (2.4)
@@
$k_{\mathrm{s}}(M)=2.5/R\quad\text{and}\quad R=\left(\frac{3M}{4\pi\bar{\rho}}\right)^{1/3}.$ (2.5)
@@
$f(\sigma)=A\left[\left(\frac{\sigma}{b}\right)^{-a}+1\right]e^{-c/\sigma^{2}},$ (2.6)
where $A$ controls the amplitude, $a$ controls the tilt, $b$ sets the mass scale at which the power law becomes significant, and $c$ determines the high-mass cutoff scale above which halo abundances exponentially decrease. This function is widely used, and it has successfully reproduced the HMFs at low z predicted by N-body simulations [40, 77]. We use galacticus [^2] https://github.com/galacticusorg/galacticus code, which utilizes the ePS formalism, to calculate the HMF for the given initial conditions generated using CLASS up to $z=500$. We choose $10^{5}M_{\odot}$ as the resolution in galacticus in order to capture the details of the smallest scales important for the Global signal. As discussed in Ref. [42], the size of the sound horizon is evolving, and due to interactions, the structure formation may be affected, especially at late times. Such interactions can suppress the growth of small-scale structures, leading to a modification in the HMF by reducing the abundance of low-mass halos compared to the standard CDM scenario.
2.1 21-cm Global Signal
@@
$\delta T_{\mathrm{b}}\simeq$ $27(1-\bar{x}_{i})\left(\frac{1-Y_{p}}{0.76}\right)\left(\frac{\Omega_{b}h^{2}}{0.023}\right)\left(\frac{0.15}{\Omega_{m}h^{2}}\frac{1+z}{10}\right)^{1/2}\left(1-\frac{T_{\mathrm{\gamma}}}{T_{\mathrm{S}}}\right)mK,$ (2.7)
where $x_{i}$ is the mean ionized fraction, $Y_{\mathrm{p}}$ is the Helium mass fraction, $T_{S}$ is the spin temperature, and $T_{\mathrm{\gamma}}$ is the CMB temperature. The spin temperature, $T_{\mathrm{S}}$, in Eq. 2.7 depends on three competing processes: coupling to the CMB photons, collisional coupling with the gas, and Lyman-$\alpha$ coupling:
$T_{\mathrm{S}}^{-1}=\frac{T_{\mathrm{\gamma}}^{-1}+x_{\mathrm{c}}T_{\mathrm{K}}^{-1}+x_{\mathrm{\alpha}}T_{\mathrm{\alpha}}^{-1}}{1+x_{\mathrm{c}}+x_{\mathrm{\alpha}}},$ (2.8)
where $x_{\mathrm{c}}$ is the collisional coupling coefficient [108], $T_{\mathrm{K}}$ is the kinetic temperature of the gas, and $x_{\mathrm{\alpha}}$ is the coefficient of Lyman-$\alpha$ coupling with $T_{\mathrm{\alpha}}$ being the corresponding Lyman-$\alpha$ color temperature. In this work, we use a modified version of ares [^3] https://github.com/treydriskell/ares/tree/galacticus_hmf [68] to predict the Global signal utilizing the output obtained from galacticus for a given set of astrophysical and IDM parameters. In the computation of the global 21-cm signal, ares takes into account for the Lyman-$\alpha$ flux, X-ray heating, and ionizing emission effects in the presence of DM-baryon interactions. Key modifications to ares are the gas kinetic temperature and the Lyman-$\alpha$ coupling [^4] For more details we refer the reader to Ref. [42].
@@
$\epsilon_{\nu}=f_{\ast}\overline{\rho}_{\mathrm{b,0}}\frac{df_{\rm{coll}}}{dt}l_{\nu},$ (2.9)
@@
where $f_{\ast}$ is the star formation efficiency, $\overline{\rho}_{b,0}$ is the mean baryon density, $f_{\rm{coll}}$ is the fraction of matter in collapsed halos with virial temperatures in excess of $T_{\mathrm{min}}$, and $l_{\nu}$ encodes the efficiency of photon production (per stellar baryon) as a function of photon frequency. Note that the product $f_{\ast}\overline{\rho}_{\mathrm{b,0}}df_{\rm{coll}}/dt$ is equivalent to the star formation rate density (SFRD) of the Universe in this model.
Figure 1: The Global 21-cm signal as a function of frequency in IDM models with $n=0$ (left-hand panel) and n=4n=-4 (right-hand panel), for DM particle mass of 11 GeV, are shown. The error bars represents the rms noise of the EDGES experiment. The color scheme captures the change in the signal as a function of the interaction cross section.
@@
We treat the $\alpha, Lyman$-continuum, and X-ray bands separately, assigning each its own distinct efficiency factor $l_{\nu}$. For the Lyman-$\alpha$ photons, we set $l_{\nu}=N_{\rm{lw}}$, where $N_{\rm{lw}}$ is the number of photons emitted per stellar baryon. For ionizing photons, $l_{\nu}=f_{\mathrm{esc}}N_{\mathrm{ion}}$. Here, we account for the possibility that only a fraction, $f_{\mathrm{esc}}$, of the ionizing photons emitted by stars, $N_{\mathrm{ion}}$, actually escape their host galaxies. Finally, for X-rays, $l_{\nu}$ is linked to the efficiency of X-ray production in nearby star-forming galaxies. The soft X-ray luminosity $L_{\mathrm{X}}$ follows the relation $L_{\mathrm{X}}\simeq 2.6\times 10^{39}\mathrm{erg}\ \rm{s}^{-1}(M_{\odot}/\rm{yr})^{-1}\left(\mathrm{SFR}/(M_{\odot}/\rm{yr})\right)$, where SFR is the star formation rate [66]. We introduce a free parameter $f_{\mathrm{X}}$ that scales this relation, i.e., $f_{\mathrm{X}}=1$ indicates the local $L_{\mathrm{X}}/\rm{SFR}$ relation, while $f_{\mathrm{X}}>1$ indicates more efficient X-ray production (per unit star formation) than $z\sim 0$ galaxies. Note that $N_{\mathrm{lw}}$, the product $f_{\mathrm{esc}}N_{\mathrm{ion}}$, and $\mathrm{f}_{\mathrm{X}}$ are all degenerate with $f_{\ast}$. Consequently, parameter forecasts are presented in terms of the product of $f_{\ast}$ and these parameters, and $f_{\ast}$ is not independently varied in this work. It is possible to break this degeneracy by independently constraining $f_{\ast}$ via high-$z$ galaxy luminosity functions [67], and performing a joint LF/21-cm likelihood analysis [54, 56]. However, we defer an exploration of this approach to future work.
@@
With the UV and X-ray emissivity in hand, one can determine the Lyman-$\alpha$ intensity as a function of redshift, as well as the ionization and heating rates of the intergalactic medium (IGM). These calculations are performed within a two-zone model of the IGM, where the fully-ionized regions and "bulk IGM" outside ionized bubbles are modeled separately. As a result, the mean ionized fraction $\bar{x}_{i}$ reflects both the volume filling factor of ionized gas, typically denoted as $Q$, and the ionized fraction of the bulk neutral IGM, $x_{e}$, such that $\bar{x}_{i}=Q+x_{e}(1-Q)$. The spin temperature represents the temperature of the largely neutral bulk IGM. For further details, see [68].
@@
We present the 21-cm Global signals computed within the IDM cosmology for various cross section values and $m_{\mathrm{\chi}}=1$ GeV in Fig. 1, along with the error bars for an EDGES-like experiment. The cosmological and astrophysical parameters are kept fixed at their fiducial values for this plot. The left-hand panel corresponds to the $n=0$ (velocity-independent cross section) model, while the right-hand panel represents the n=4n=-4 (Coulomb-like interaction) model. From the left-hand panel, it is evident that for the velocity-independent cross section model, the signals show a smooth shift of the troughs towards higher frequencies as the cross section increases. However, in case of Coulomb-like interaction, the signal exhibits a non-monotonic behavior in the position of the troughs with increasing cross sections. This difference in the behavior of the 21-cm signal with cross section for two IDM models can be attributed to the competition of two effects: the dissipation of the relative bulk velocity of DM and baryons and the impact of interaction cross section on the efficiency of baryon cooling.
3 Methods
3.1 Fisher Forecasting
$F_{\mathrm{ij}}=-\left<\frac{\partial^{2}\ln\mathcal{L}}{\partial\theta_{\mathrm{i}}\partial\theta_{\mathrm{j}}}\right$>\,, (3.1)
where $F_{\mathrm{ij}}$ are the $(\mathrm{i},\mathrm{j})$-th elements of the Fisher matrix, $\theta_{\mathrm{i}}$ are the model parameters, $\ln\mathcal{L}$ represents the log-likelihood function, and <><> denotes the ensemble average over different data realizations.
$C_{\mathrm{ij}}=F^{-1}_{\mathrm{ij}}\,$, (3.2)
where $C_{\mathrm{ij}}$ are the $(\mathrm{i},\mathrm{j})$-th elements of the covariance matrix and $F^{-1}_{\mathrm{ij}}$ are the $(\mathrm{i},\mathrm{j})$-th element of the inverse of Fisher matrix. Once the Fisher matrix is calculated, one can compute the uncertainties (standard deviations) of individual parameters $\theta_{\mathrm{i}}$ as follows:
$\sigma(\theta_{\mathrm{i}})=\sqrt{F^{-1}_{\mathrm{ii}}}\,$, (3.3)
$\rho_{\mathrm{ij}}=\frac{C_{\mathrm{ij}}}{\sigma(\theta_{\mathrm{i}})\sigma(\theta_{\mathrm{j}})}\,.$ (3.4)
$\mathcal{L}(\boldsymbol{\mu}|\boldsymbol{X})=\frac{\exp\Bigl(-\frac{1}{2}(\boldsymbol{X}-\boldsymbol{\mu})^{\mathrm{T}}\boldsymbol{\Sigma}^{-1}(\boldsymbol{X}-\boldsymbol{\mu})\Bigl)}{\sqrt{(2\pi)^{\mathrm{k}}|\boldsymbol{\Sigma}|}}\,$, (3.5)
where $\boldsymbol{X}$ is the observed $k$-dimensional column vector, $\boldsymbol{\mu}$ is the model-predicted $k$-dimensional column vector, $\boldsymbol{\Sigma}$ is the symmetric input covariance matrix, and $|\boldsymbol{\Sigma}|$ is the determinant of the input covariance matrix.
In the case of 21-cm cosmology, the log-likelihood function is
$\ln\mathcal{L}(\boldsymbol{\theta}|\boldsymbol{X})\propto-\frac{1}{2}\Bigl((\boldsymbol{X}-T(\boldsymbol{\theta}))^{T}\boldsymbol{\Sigma}^{-1}(\boldsymbol{X}-T(\boldsymbol{\theta}))\Bigl)\,$, (3.6)
where $\boldsymbol{\theta}$ is the parameter column vector and $T(\boldsymbol{\theta})$ is the temperature of the global 21-cm signal generated by ares code for $\boldsymbol{\theta}$ as its input parameters. In the following subsection, we will elaborate more on the models used for the input covariance matrix.
Experiment $\nu_{\mathrm{min}},\nu_{\mathrm{max}}$ (MHz) $N_{\mathrm{ch}}$ $\Delta\nu$ (kHz) $\tau_{\mathrm{obs}}$ (h) $T_{\mathrm{rec}}$ (K) $\sigma_{\mathrm{RMS}}\ (\mathrm{mK})$ Input Noise